### v2

@@ 2.5 Estimation and Statistical Inference
$\left[\hat{\delta}_{ATE}-1.96\sqrt{\operatorname{Var}\left(\hat{\delta}_{ATE}\right)},\hat{\delta}_{ATE}+1.96\sqrt{\operatorname{Var}\left(\hat{\delta}_{ATE}\right)}\right],.$
However, without empirical information of $\operatorname{Cov}\left(\hat{\gamma}_{1},\hat{\alpha}_{1}\right),\operatorname{Cov}\left(\hat{\gamma}_{2},\hat{\alpha}_{1}\right)$, and $\operatorname{Cov}\left(\hat{\gamma}_{3},\hat{\alpha}_{1}\right)$, the analyst may ignore the estimation uncertainty in the Stage 1 analysis by viewing $\widehat{\alpha}_{1}$ as a constant. Additionally, because the sample estimates $\hat{\alpha}_{1k},\hat{\beta}_{1k}$, and $\hat{\beta}_{2k}$ obtained from the Stage 1 analysis are used as predictors in the Stage 2 analysis, estimators of the coefficients in the Stage 2 regression $\hat{\gamma}_{1},\hat{\gamma}_{2},\hat{\gamma}_{3}$, and $\hat{\theta}_{V}$ and their variances and covariances likely contain bias. These issues potentially jeopardize the statistical inference for $\delta_{\text{ATE }}$ (Cameron and Trivedi, 2005).
Multilevel bootstrap BCa. To account for Stage 1 estimation uncertainty in two-stage analyses, researchers often resort to bootstrapping for robust inference. In this study, we adopt a multilevel bootstrapping approach (Goldstein, 2011). We first generate a bootstrap sample of sites through a simple random resampling with replacement of the site indices. We then generate a bootstrap sample of individuals within each treatment group at each bootstrapped site. By repeating this procedure, we generate a total of $B$ bootstrapped samples and obtain $B$ bootstrapped estimates of $\delta_{ATE}$. Rather than using the conventional percentile method to construct the $95\%$ confidence interval, we opt for the bias-corrected accelerated (BCa) method. The latter adjusts for both bias and skewness in the bootstrap distribution and allows the standard error to be a function of the parameter value. Therefore, BCa is second-order accurate and is preferrable to the percentile method that is only first-order accurate (Efron and Tibshirani, 1993).
@@ 3.1 Research Questions for the Simulations
We empirically address the following two questions through simulations. The first question is how the proposed MS2T-IV strategy compares to existing alternative methods in estimating the cumulative ATE; the second question is how the coverage rates compare between the improper inference strategy and the multiple bootstrap BCa strategy.
@@ 3.2 Data Generation
We generate simulation data in accordance with the theoretical models (1). Baseline covariates include an observed $X$ and an unobserved $U$. Following typical multisite designs in application research, we specify the number of sites to be either 25 or 100 and let the sample size per site vary from 30 to 5,000. We generate 500 independent data sets in each case. Details of the data generation are in section S3 of the supplementary material.
@@ 3.3 Simulation Results: Estimation
Efficiency. Across all the methods under consideration, the empirical variance decreases when either the number of sites or the sample size per site increases. The naive estimator and the IPTW estimator show comparable levels of efficiency, while the MS2T-IV estimator is relatively less efficient. For the naive estimator and the MS2T-IV estimator, covariance adjustment for $X$ improves efficiency to a limited extent.
MSE. Among all the estimators compared, only the MS2T-IV estimators are consistent and show the fastest rate of MSE reduction when the number of sites or the sample size per site increases. In contrast, because the naive estimators and the IPTW estimator are biased, their MSE does not converge to zero, even with an increase in the number of sites or the sample size per site.
@@
Naive 4.82 4.90 4.84 4.80 4.77 4.87 4.83 4.86
Naive-adj 3.78 3.95 3.94 3.89 3.80 3.95 3.92 3.94
IPTW 4.69 3.15 2.39 2.34 4.65 3.10 2.37 2.39
MS2T-IV 1.96 1.29 0.22 0.00 2.26 1.37 0.21 0.14
MS2T-IV-adj 1.45 1.13 0.20 0.01 1.73 1.09 0.16 0.12
@@
Naive 6.16 2.29 0.62 0.50 1.56 0.56 0.15 0.13
Naive-adj 6.41 2.03 0.55 0.48 1.63 0.48 0.14 0.12
IPTW 6.65 2.24 0.58 0.46 1.71 0.55 0.14 0.12
MS2T-IV 37.8 26.8 10.3 7.72 9.73 5.33 2.32 1.89
MS2T-IV-adj 34.5 23.0 9.76 7.54 8.23 4.89 2.08 1.84
@@
Naive 29.4 26.3 24.1 23.6 24.3 24.2 23.5 23.7
Naive-adj 20.7 17.6 16.1 15.6 16.1 16.1 15.5 15.6
IPTW 28.7 12.2 6.29 5.94 23.3 10.2 5.78 5.85
MS2T-IV 41.6 28.5 10.4 7.72 14.8 7.20 2.36 1.91
MS2T-IV-adj 36.6 24.3 9.80 7.54 11.2 6.09 2.10 1.86
@@ 4 Application Study
Our causal estimand is the cumulative ATE of attending a small class versus a regular class over the two consecutive years. The target population for this study is represented by the 4,511 students who were enrolled in 76 Project STAR schools in both kindergarten and Grade 1. However, noncompliance started to occur at the beginning of Grade 1: 108 students switched from a small class to a regular class, while 248 students switched from a regular class to a small class at the end of the kindergarten year. The noncompliance rate was about $8\%$ in each treatment group. Krueger (1999) speculated that these moves between class size types were likely due to behavioral problems or parental complaints.
The outcome measure is the sum of reading and math test scores by the end of Grade 1. We also measure the sum of reading and math test scores by the end of kindergarten, which is a key posttreatment confounder in this study. Table 4 compares the actual class sizes and achievement scores in kindergarten and Grade 1 across the four treatment sequences. A naive comparison between students who stayed in a small class and those who stayed in a regular class over both kindergarten and Grade 1 shows a 24-point difference in Grade 1 achievement. Yet notably, students who switched from a small class to a regular class ($Z=1$, $D=0$) displayed significantly lower kindergarten scores on average compared to those who remained in a small class $(Z=1,D=1)$ : 919 vs. 943. Similarly, students who switched from a regular class to a small class ($Z=0,D=1$) displayed lower kindergarten scores than those who remained in a regular class ($Z=0,D=0$): 921 vs. 928. The between-group difference in the former case is about three times the magnitude of the latter. Hence the naive comparison between the ($Z=1,D=1)$ group and the ($Z=0,D=0$) group would likely contain a positive bias. Further analysis reveals that the sum of kindergarten reading and math scores is a statistically significant predictor of compliance among those in the $Z=1$ group only.
We further observe that, in both the $Z=0$ group and $Z=1$ group, the noncompliance rate is considerably higher among students whose within-school relative standings fall within the bottom two quintiles of kindergarten scores. Based on this, we create a binary measure of $V$, using the within-school 40th percentile of kindergarten scores as the cutoff.
@@
According to our analytic results aggregated over the ten imputed datasets, attending a small class as opposed to a regular class in both kindergarten and Grade 1 is expected to increase the sum of the Grade 1 reading and math scores by 21.03 points. Using the within-school standard deviation of the outcome in the control group as the scaling unit, the effect size is about 0.27. The improper $95\%$ confidence interval for the cumulative ATE is $(6.80,35.25)$ while the Bootstrap BCa confidence interval is ($3.72,69.08$). In the latter case, we generated 500 bootstrapped samples. For each bootstrapped sample, we calculated the average estimate of the cumulative ATE across the 10 imputed datasets. Based on the 500 bootstrapped estimates, we constructed the Bootstrap BCa confidence interval. We find that the conclusion is insensitive to the choice of the cutoff value for the kindergarten scores.
@@ 5.1 Contributions of this study
According to the simulation results, in comparison with the naive method that adjusts only for observed pretreatment covariates and with the IPTW method that adjusts for observed pretreatment and posttreatment covariates, the MS2T-IV method excels in bias reduction. Although the MS2T-IV estimator is generally less efficient than these existing alternatives, it outperforms them in MSE reduction when the number of sites is sufficient large.
@@ 5.2 Limitations and Future Research
Second, as suggested by our simulation results, adjusting for site mean-centered pretreatment covariates may help reduce hidden bias associated with unobserved posttreatment confounders, to the extent that such confounders are correlated with the observed pretreatment covariates. This adjustment enhances the plausibility of the within-site zero covariance assumption. Additionally, it reduces finite-sample bias and improves estimation efficiency. Future research may further quantify the remaining finite-sample bias and explore strategies for bias correction.
Third, like the MSMM-IV method, the MS2T-IV method relies on the assumption of between-site independence. This assumption can be challenged on scientific grounds but cannot be empirically verified. Analysts may include observed site-level pretreatment covariates in the Stage 2 analysis to help mitigate potential bias. Future research may also explore sensitivity analysis strategies to assess potential consequences of its violations.
Fourth, it is not uncommon that a treatment sequence may consist of more than two phases. Future research may extend the proposed strategy to evaluations of the cumulative ATE of multi-phase interventions.
@@ 8 Acknowledgments
To simplify the presentation, we define $\alpha_{hik}^{\prime}=\alpha_{hik}-\alpha_{hk}$ for $h=0,1;\beta_{Dhik}^{\prime}=\beta_{Dhik}-\beta_{Dhk}$ for $h=0,1;\theta_{Y0ik}^{\prime}=\theta_{Y0ik}-\theta_{Y0k};and\gamma_{hik}^{\prime}=\gamma_{hik}-\gamma_{hk}$ for $h=1,2,3.$
S1.1 Reduced-Form Model for V
Here we derive the results for the reduced-form model for $V$. For each school $k$, following the theoretical model for $V_{ik}(z)$, we have that $E\left\{V_{ik}(z)\mid S_{i}=k\right\}=E\left(\alpha_{0ik}+\alpha_{1ik}z\mid S_{i}=k\right)=\alpha_{0k}+\alpha_{1k}z$, and therefore, $E\left\{V_{ik}(1)-V_{ik}(0)\mid S_{i}=k\right\}=\alpha_{1k}$. The analytic model for $V$ in school $k$ is
$V_{ik}$ $=\left(1-Z_{ik}\right)V_{ik}(0)+Z_{ik}V_{ik}(1)$
$=\left(1-Z_{ik}\right)\alpha_{0ik}+Z_{ik}\left(\alpha_{0ik}+\alpha_{1ik}\right)$
$=\alpha_{0ik}+\alpha_{1ik}Z_{ik},$
Then
$V_{ik}=\alpha_{0k}+\alpha_{1k}Z_{ik}+e_{Vik},$
where $e_{Vik}=\alpha_{0ik}^{\prime}+\alpha_{1ik}^{\prime}Z_{ik}$. Since $Z_{ik}$ is randomized, we can prove that $e_{Vik}$ has mean zero and is uncorrelated with $Z_{ik}$ within school $k$.
$E\left(e_{Vik}\mid S_{i}=k\right)=E\left(\alpha_{0ik}^{\prime}\mid S_{i}=k\right)+E\left(\alpha_{1ik}^{\prime}Z_{ik}\mid S_{i}=k\right)=0.$
Further note that $Z_{ik}\cdot Z_{ik}=Z_{ik}$, and thus
$E\left(e_{Vik}Z_{ik}\mid S_{i}=k\right)=E\left(\alpha_{0ik}^{\prime}Z_{ik}\mid S_{i}=k\right)+E\left(\alpha_{1ik}^{\prime}Z_{ik}\mid S_{i}=k\right)=0$
S1.2 Reduced-Form Model for D
Here we derive the results for the reduced-form model for $D$. For each school $k$, following the theoretical model for $D_{ik}(z)$, we have that
$E\left\{D_{ik}(z)\mid S_{i}=k\right\}$ $=E\left\{\beta_{D0ik}+\beta_{D1ik}z+\beta_{Vzk}V_{ik}(z)+e_{z.ik}^{(D)}\mid S_{i}=k\right\}$
$=\beta_{D0k}+\beta_{D1k}z+\beta_{Vzk}E\left\{V_{ik}(z)\mid S_{i}=k\right\}$
$=\beta_{D0k}+\beta_{D1k}z+\beta_{Vzk}\left(\alpha_{0k}+\alpha_{1k}z\right),$
where $\beta_{Dhk}=E\left(\beta_{Dhik}\mid S_{i}=k\right)$ for $h=0,1$. Let
$\begin{gathered}\beta_{0k}=\beta_{D0k}+\alpha_{0k}\beta_{V0k};\\ \beta_{1k}=\beta_{D1k}+\left(\alpha_{0k}+\alpha_{1k}\right)\beta_{V1k}-\alpha_{0k}\beta_{V0k}.\end{gathered}$
Then, $E\left\{D_{ik}(0)\mid S_{i}=k\right\}=\beta_{0k}$, $E\left\{D_{ik}(1)\mid S_{i}=k\right\}=\beta_{0k}+\beta_{1k}$, and $E\left\{D_{ik}(1)-D_{ik}(0)\mid S_{i}=k\right\}=\beta_{1k}$.
The analytic model for $D$ in school $k$ is
$D_{ik}=\left(1-Z_{ik}\right)D_{ik}(0)+Z_{ik}D_{ik}(1)$
$\left(1-Z_{ik}\right)\left\{\beta_{D0ik}+\beta_{V0k}V_{ik}(0)+e_{0.ik}^{(D)}\right\}+Z_{ik}\left\{\beta_{D0ik}+\beta_{D1ik}+\beta_{V1k}V_{ik}(1)+e_{1.ik}^{(D)}\right\}$
$\beta_{D0ik}+\beta_{D1ik}Z_{ik}+\beta_{V0k}\alpha_{0ik}\left(1-Z_{ik}\right)+\beta_{V1k}\left(\alpha_{0ik}+\alpha_{1ik}\right)Z_{ik}+e_{ik}^{(D)},$
where $e_{ik}^{(D)}=\left(1-Z_{ik}\right)e_{0.ik}^{(D)}+Z_{ik}e_{1.ik}^{(D)}$. Then
$D_{ik}=\beta_{D0k}+\alpha_{0k}\beta_{V0k}+\left\{\beta_{D1k}+\left(\alpha_{0k}+\alpha_{1k}\right)\beta_{V1k}-\alpha_{0k}\beta_{V0k}\right\}Z_{ik}+e_{Dik},$
where $e_{Dik}=\beta_{D0ik}^{\prime}+\alpha_{0\mathrm{i}k}^{\prime}\beta_{V0k}+\left\{\beta_{D1ik}^{\prime}+\beta_{V1k}\left(\alpha_{0\mathrm{i}k}^{\prime}+\alpha_{1\mathrm{i}k}^{\prime}\right)-\beta_{V0k}\alpha_{0\mathrm{i}k}^{\prime}\right\}Z_{ik}+e_{ik}^{(D)}$. Then
$D_{ik}=\beta_{0k}+\beta_{1k}Z_{ik}+e_{Dik}.$
Given that $Z_{ik}$ is randomized, we can show that $E\left(e_{Dik}\mid S_{i}=k\right)=E\left(e_{Dik}Z_{ik}\mid S_{i}=k\right)=0$
$~~~~~E\left(e_{Dik}\mid S_{i}=k\right)$
$=E\left(\beta_{D0ik}^{\prime}\mid S_{i}=k\right)+E\left(\alpha_{0ik}^{\prime}\beta_{V0k}\mid S_{i}=k\right)$
$+E\left[\left\{\beta_{D1ik}^{\prime}+\beta_{V1k}\left(\alpha_{0ik}^{\prime}+\alpha_{1ik}^{\prime}\right)-\beta_{V0k}\alpha_{0ik}^{\prime}\right\}Z_{ik}\mid S_{i}=k\right]+E\left(e_{ik}^{(D)}\mid S_{i}=k\right)$
$=0;$
$~~~~~E\left(e_{Dik}Z_{ik}\mid S_{i}=k\right)$
$=E\left(\beta_{D0ik}^{\prime}Z_{ik}\mid S_{i}=k\right)+E\left(\alpha_{0ik}^{\prime}\beta_{V0k}Z_{ik}\mid S_{i}=k\right)$
$+E\left[\left\{\beta_{D1ik}^{\prime}+\beta_{V1k}\left(\alpha_{0ik}^{\prime}+\alpha_{1ik}^{\prime}\right)-\beta_{V0k}\alpha_{0ik}^{\prime}\right\}Z_{ik}\mid S_{i}=k\right]+E\left\{e_{ik}^{(D)}Z_{ik}\mid S_{i}=k\right\}$
$=0;$
S1.3 Reduced-Form Model for $\boldsymbol{Y}$
@@
$=E\left\{\theta_{Y0ik}+\theta_{Vk}V_{ik}(0)+\gamma_{2ik}D_{ik}(0)\mid S_{i}=k\right\}$
$=\theta_{Y0k}+\theta_{Vk}\alpha_{0k}+E\left\{\gamma_{2ik}\left(\beta_{D0ik}+\beta_{V0k}V_{ik}(0)+e_{0.ik}^{(D)}\right)\mid S_{i}=k\right\}$
$=\theta_{Y0k}+\theta_{Vk}\alpha_{0k}+E\left\{\gamma_{2ik}\left(\beta_{D0ik}+\beta_{V0k}\alpha_{0ik}+e_{0.ik}^{(D)}\right)\mid S_{i}=k\right\}.$
Under Assumption 2, $\operatorname{Cov}\left(\gamma_{2ik},\beta_{D0ik}\mid S_{i}=k\right)=\operatorname{Cov}\left(\gamma_{2ik},\alpha_{0ik}\mid S_{i}=k\right)=0$, we have that
$\begin{gathered}E\left\{Y_{ik}(0)\mid S_{i}=k\right\}=\theta_{Y0k}+\theta_{Vk}\alpha_{0k}+\gamma_{2k}\beta_{0k},\\ E\left\{Y_{ik}(1)\mid S_{i}=k\right\}=E\left[Y_{ik}\left(1,D_{ik}(1)\right)\mid S_{i}=k\right]\\ =E\left\{\theta_{Y0ik}+\theta_{Vk}V_{ik}(1)+\gamma_{1ik}+\gamma_{2ik}D_{ik}(1)+\gamma_{3ik}D_{ik}(1)\mid S_{i}=k\right\}\\ =\theta_{Y0k}+\theta_{Vk}\left(\alpha_{0k}+\alpha_{1k}\right)+\gamma_{1k}+E\left[(\gamma_{2ik}+\gamma_{3ik})\left\{\beta_{D0ik}+\beta_{D1ik}\right.\right.\\ \left.\left.+\beta_{V1k}\left(\alpha_{0ik}+\alpha_{1ik}\right)+e_{z.ik}^{(D)}\right\}\mid S_{i}=k\right]\\ =\theta_{Y0k}+\theta_{Vk}\left(\alpha_{0k}+\alpha_{1k}\right)+\gamma_{1k}+\left(\gamma_{2k}+\gamma_{3k}\right)\left(\beta_{0k}+\beta_{1k}\right),\end{gathered}$
where the last identity holds under Assumption 2 that $\operatorname{Cov}\left(\gamma_{2ik},\beta_{D0ik}\mid S_{i}=k\right)=$
$\operatorname{Cov}\left(\gamma_{2ik},\alpha_{0ik}\mid S_{i}=k\right)=\operatorname{Cov}\left(\gamma_{2ik},\beta_{D1ik}\mid S_{i}=k\right)=\operatorname{Cov}\left(\gamma_{2ik},\alpha_{1ik}\mid S_{i}=k\right)=$
$\operatorname{Cov}\left(\gamma_{3ik},\beta_{D0ik}\mid S_{i}=k\right)=\operatorname{Cov}\left(\gamma_{3ik},\alpha_{0ik}\mid S_{i}=k\right)=\operatorname{Cov}\left(\gamma_{3ik},\beta_{D1ik}\mid S_{i}=k\right)=\operatorname{Cov}\left(\gamma_{3ik},\alpha_{1ik}\mid S_{i}=k\right)=0$.
The ITT effect of $Z$ on $Y$ at site k is
where $\beta_{2k}=\beta_{0k}+\beta_{1k}$.
The analytic model for obtaining the ITT effect on $Y$ at site $k$ is
$Y_{ik}=\left(1-Z_{ik}\right)Y_{ik}\left(0,D_{ik}(0)\right)+Z_{ik}Y_{ik}\left(1,D_{ik}(1)\right)$
$=\left(1-Z_{ik}\right)\left\{\theta_{Y0ik}+\theta_{Vk}V_{ik}(0)+\gamma_{2ik}D_{ik}(0)\right\}+Z_{ik}\left\{\theta_{Y0ik}+\theta_{Vk}V_{ik}(1)+\gamma_{1ik}+\gamma_{2ik}D_{ik}(1)\right$.
$\left.\quad+\gamma_{3ik}D_{ik}(1)\right\}$
$=\theta_{Y0ik}+\theta_{Vk}V_{ik}(0)+\theta_{Vk}\left\{V_{ik}(1)-V_{ik}(0)\right\}Z_{ik}+\gamma_{1ik}Z_{ik}$
$\quad+\gamma_{2ik}\left\{\left(1-Z_{ik}\right)D_{ik}(0)+Z_{ik}D_{ik}(1)\right\}+\gamma_{3ik}Z_{ik}D_{ik}(1)$
$=\theta_{Y0ik}+\theta_{Vk}\alpha_{0ik}+\theta_{Vk}\alpha_{1ik}Z_{ik}+\left(\gamma_{1k}+\gamma_{1ik}^{\prime}\right)Z_{ik}+\left(\gamma_{2k}+\gamma_{2ik}^{\prime}\right)D_{ik}+\left(\gamma_{3k}+\gamma_{3ik}^{\prime}\right)D_{ik}Z_{ik}$
$=\theta_{Y0ik}+\theta_{Vk}\alpha_{0ik}+\theta_{Vk}\alpha_{1ik}Z_{ik}+\left(\gamma_{1k}+\gamma_{1ik}^{\prime}\right)Z_{ik}+\left(\gamma_{2k}+\gamma_{2ik}^{\prime}\right)\left(\beta_{0k}+\beta_{1k}Z_{ik}+e_{Dik}\right)$
$\quad+\left(\gamma_{3k}+\gamma_{3ik}^{\prime}\right)\left(\beta_{0k}Z_{ik}+\beta_{1k}Z_{ik}+e_{Dik}Z_{ik}\right)$
$=\left(\theta_{Y0k}+\theta_{Vk}\alpha_{0k}+\gamma_{2k}\beta_{0k}\right)+\left\{\theta_{Vk}\alpha_{1k}+\gamma_{1k}+\gamma_{2k}\beta_{1k}+\gamma_{3k}\left(\beta_{0k}+\beta_{1k}\right)\right\}Z_{ik}+\left\{\theta_{Y0ik}^{\prime}\right$.
$\quad+\theta_{Vk}\alpha_{0ik}^{\prime}+\theta_{Vk}\alpha_{1ik}^{\prime}Z_{ik}+\gamma_{1ik}^{\prime}Z_{ik}+\gamma_{2k}e_{Dik}+\gamma_{2ik}^{\prime}\left(\beta_{0k}+\beta_{1k}Z_{ik}+e_{Dik}\right)$
$\left.\quad+\gamma_{3k}e_{Dik}Z_{ik}+\gamma_{3ik}^{\prime}\left(\beta_{0k}Z_{ik}+\beta_{1k}Z_{ik}+e_{Dik}Z_{ik}\right)\right\}$
$=\theta_{0k}+\theta_{1k}Z_{ik}+e_{Yik}$
@@
Here, $\theta_{0k}=\theta_{\mathrm{Y}0k}+\theta_{Vk}\alpha_{0k}+\gamma_{2k}\beta_{0k};$
@@
$\theta_{1k}=\theta_{Vk}\alpha_{1k}+\gamma_{1k}+\gamma_{2k}\beta_{1k}+\gamma_{3k}\left(\beta_{0k}+\beta_{1k}\right);$

### v3

@@ 2.6 Estimation and Statistical Inference
$\left[\hat{\delta}_{ATE}-1.96\sqrt{\operatorname{Var}\left(\hat{\delta}_{ATE}\right)},\hat{\delta}_{ATE}+1.96\sqrt{\operatorname{Var}\left(\hat{\delta}_{ATE}\right)}\right].$
However, without empirical information of $\operatorname{Cov}\left(\hat{\gamma}_{1},\hat{\alpha}_{1}\right),\operatorname{Cov}\left(\hat{\gamma}_{2},\hat{\alpha}_{1}\right)$, and $\operatorname{Cov}\left(\hat{\gamma}_{3},\hat{\alpha}_{1}\right)$, the analyst may ignore the estimation uncertainty in the Stage 1 analysis by viewing $\widehat{\alpha}_{1}$ as a constant. Additionally, because the sample estimates $\hat{\alpha}_{1k},\hat{\beta}_{1k}$, and $\hat{\beta}_{2k}$ obtained from the Stage 1 analysis are used as predictors in the Stage 2 analysis, estimators of the coefficients in the Stage 2 regression $\hat{\gamma}_{1},\hat{\gamma}_{2},\hat{\gamma}_{3}$, and $\hat{\theta}_{V}$ and their variances and covariances likely contain bias. These issues potentially jeopardize the statistical inference for $\delta_{\text{ATE }}$ (CameronTrivedi2005).
Multilevel bootstrap BCa. To account for Stage 1 estimation uncertainty in two-stage analyses, researchers often resort to bootstrapping for robust inference. In this study, we adopt a multilevel bootstrapping approach (Goldstein2011). We first generate a bootstrap sample of sites through a simple random resampling with replacement of the site indices. We then generate a bootstrap sample of individuals within each treatment group at each bootstrapped site. By repeating this procedure, we generate a total of $B$ bootstrapped samples and obtain $B$ bootstrapped estimates of $\delta_{ATE}$. Rather than using the conventional percentile method to construct the $95\%$ confidence interval, we opt for the bias-corrected accelerated (BCa) method. The latter adjusts for both bias and skewness in the bootstrap distribution and allows the standard error to be a function of the parameter value. Therefore, BCa is second-order accurate and is preferrable to the percentile method that is only first-order accurate (EfronTibshirani1993).
@@ 3.1 Research Questions for the Simulations
We empirically address the following two questions through simulations. The first question is how the proposed MS2T-IV strategy compares to existing alternative methods in estimating the cumulative ATE; the second question is how the coverage rates compare between the improper inference strategy and the multiple bootstrap BCa strategy. In addition, we examine through simulations the consequences when Assumption 2 and 3 are violated.
@@ 3.2 Data Generation
We generate simulation data in accordance with the theoretical model (1). Baseline covariates include an observed $X$ and an unobserved $U$. Following typical multisite designs in application research, we specify the number of sites to be either 25 or 100 and let the sample size per site vary from 30 to 5,000. We generate 500 independent data sets in each case. Details of the data generation are in section S3 of the supplementary material.
@@ 3.3 Simulation Results: Estimation
Efficiency. Across all the methods under consideration, the empirical variance decreases when either the number of sites or the sample size per site increases. The MS2T-IV estimator is relatively less efficient than the naive estimator but relatively more efficient than the IPTW estimator. For the naive estimator and the MS2T-IV estimator, covariance adjustment for $X$ improves efficiency to a limited extent.
MSE. Among all the estimators compared, only the MS2T-IV estimators are consistent and show a fast rate of MSE reduction when the number of sites or the sample size per site increases. In contrast, because the naive estimators and the IPTW estimator are biased, their MSE does not converge to zero.
When the within-site zero covariance assumption (Assumption 2) is violated, none of the five estimators will be consistent. The bias does not converge to zero as the sample size per site increases. When the between-site independence assumption (Assumption 3) is violated, the clear downward trend in bias also disappears for the MS2T-IV estimator. Nevertheless, in these cases, the MS2T-IV estimator with adjustment for $X$ still almost always has the smallest bias and MSE among all five estimators considered in this simulation study. Details of the simulation results with violated assumptions are in section S5 of the supplementary material.
@@
Naive 1.79 1.75 1.80 1.80 1.73 1.74 1.79 1.75
Naive-adj 1.35 1.32 1.38 1.38 1.23 1.33 1.38 1.33
IPTW -2.52 -1.45 -0.50 -0.37 -2.69 -1.45 -0.58 -0.49
MS2T-IV -0.05 -0.03 0.03 -0.01 -0.06 -0.03 0.03 0.01
MS2T-IV-adj -0.09 -0.07 0.02 -0.01 -0.10 -0.07 0.01 0.01
@@
Naive 5.34 2.72 0.97 0.91 1.44 0.58 0.25 0.24
Naive-adj 4.86 2.35 0.77 0.75 1.42 0.51 0.21 0.19
IPTW 25.33 9.04 4.93 4.32 6.73 2.24 1.27 1.24
MS2T-IV 8.90 4.29 1.69 1.64 2.08 1.15 0.44 0.31
MS2T-IV-adj 8.13 3.58 1.62 1.61 1.94 1.01 0.41 0.31
@@
Naive 8.53 5.78 4.20 4.15 4.43 3.60 3.46 3.29
Naive-adj 6.68 4.09 2.66 2.65 2.94 2.28 2.10 1.96
IPTW 31.67 11.13 5.18 4.45 13.96 4.36 1.60 1.48
MS2T-IV 8.91 4.30 1.70 1.64 2.08 1.15 0.44 0.31
MS2T-IV-adj 8.13 3.58 1.62 1.61 1.95 1.02 0.41 0.31
@@ 4 Application Study
Our causal estimand is the cumulative ATE of attending a small class versus a regular class over the two consecutive years. The target population for this study is represented by the 4,511 students who were enrolled in 76 Project STAR schools in both kindergarten and Grade 1. However, noncompliance started to occur at the beginning of Grade 1: 108 students switched from a small class to a regular class, while 248 students switched from a regular class to a small class at the end of the kindergarten year. The noncompliance rate was about $8\%$ in each treatment group. Krueger1999 speculated that these moves between class size types were likely due to behavioral problems or parental complaints.
The outcome measure is the sum of reading and math test scores by the end of Grade 1, while the sum of reading and math test scores by the end of kindergarten is a key posttreatment confounder in this study. Table 4 compares the actual class sizes and achievement scores in kindergarten and Grade 1 across the four treatment sequences. A naive comparison between students who stayed in a small class and those who stayed in a regular class over both kindergarten and Grade 1 shows a 24-point difference in Grade 1 achievement. Yet based on an analysis regressing the kindergarten score on the $Z$-by-$D$ group dummies with school fixed effects, students who switched from a small class to a regular class ($Z=1$, $D=0$) displayed significantly lower kindergarten scores on average compared to those who remained in a small class $(Z=1,D=1)$. Similarly, students who switched from a regular class to a small class ($Z=0,D=1$) displayed lower kindergarten scores than those who remained in a regular class ($Z=0,D=0$). The average between-group difference is 23.5-23.5 with a p-value less than 0.001 in the former case, and 1.6-1.6 with a p-value of 0.345 in the latter case. Hence the naive comparison between the ($Z=1,D=1)$ group and the ($Z=0,D=0$) group would likely contain a positive bias. Further analysis reveals that the sum of kindergarten reading and math scores is a statistically significant predictor of compliance among those in the $Z=1$ group only.
@@
According to our analytic results aggregated over the ten imputed datasets, attending a small class as opposed to a regular class in both kindergarten and Grade 1 is expected to increase the sum of the Grade 1 reading and math scores by 20.21 points. Using the within-school standard deviation of the outcome in the control group as the scaling unit, the effect size is about 0.26. The improper $95\%$ confidence interval for the cumulative ATE is $(6.91,33.51)$ while the Bootstrap BCa confidence interval is $(5.67,77.39)$. In the latter case, we generated 500 bootstrapped samples. For each bootstrapped sample, we calculated the average estimate of the cumulative ATE across the 10 imputed datasets. Based on the 500 bootstrapped estimates, we constructed the Bootstrap BCa confidence interval.
@@ 5.1 Contributions of this study
According to the simulation results, in comparison with the naive method that adjusts only for observed pretreatment covariates and with the IPTW method that adjusts for observed pretreatment and posttreatment covariates, the MS2T-IV method excels in bias reduction. Although the MS2T-IV estimator is generally less efficient than these existing alternatives, it outperforms them in MSE reduction when the number of sites is sufficiently large.
@@ 5.2 Limitations and Future Research
Second, as suggested by our simulation results, adjusting for site mean-centered pretreatment covariates may reduce finite-sample bias and improve estimation efficiency. Future research may further quantify the remaining finite-sample bias and explore strategies for bias correction.
Third, our assumptions 2 and 3 are violated if there are individual-level or site-level modifiers. So if the coefficients differ by subpopulations of individuals or subpopulations of sites, then these assumptions are likely violated. In such cases, the analysis will need to be conducted within each subpopulation. And eventually, the cumulative ATE will need to be an average over the subpopulations. Yet even if all the modifiers are observed, this may be impractical due to sample size constraints.
Fourth, it is not uncommon that a treatment sequence may consist of more than two phases. Noncompliance may occur in multiple phases including phase I. Future research may extend the proposed strategy to evaluations of the cumulative ATE of multi-phase interventions.
@@ 8 Acknowledgments
S1.1 Reduced-Form Model for $\boldsymbol{Y}$
@@
$=E\left\{\theta_{Y0ik}+\theta_{Vik}V_{ik}(0)+\gamma_{2ik}D_{ik}(0)\mid S_{i}=k\right\},$
and
$E\left\{Y_{ik}(1)\mid S_{i}=k\right\}$ $=E\left[Y_{ik}\left(1,D_{ik}(1)\right)\mid S_{i}=k\right]$
$=E\left\{\theta_{Y0ik}+\theta_{Vik}V_{ik}(1)+\gamma_{1ik}+\gamma_{2ik}D_{ik}(1)+\gamma_{3ik}D_{ik}(1)\mid S_{i}=k\right\}.$
Then, the ITT effect of $Z$ on $Y$ at site k is
$~~~~E\left\{Y_{ik}(1)-Y_{ik}(0)\mid S_{i}=k\right\}$
$=E\left[\theta_{Vik}\left\{V_{ik}(1)-V_{ik}(0)\right\}+\gamma_{1ik}+\gamma_{2ik}\left\{D_{ik}(1)-D_{ik}(0)\right\}+\gamma_{3ik}D_{ik}(1)\mid S_{i}=k\right]$
$=\theta_{Vk}\alpha_{1k}+\gamma_{1k}+\gamma_{2k}\beta_{1k}+\gamma_{3k}\beta_{2k},$
@@
under the assumption 2 that $\mathrm{Cov}\{\gamma_{2ik},\beta_{1ik}\}=\mathrm{Cov}\{\gamma_{3ik},\beta_{2ik}\}=\mathrm{Cov}\{\theta_{Vik},\alpha_{1ik}\}$.
Given that $Z_{ik}$ is randomized, the analytic model for obtaining the ITT effect on $Y$ at site $k$ is
$Y_{ik}=\theta_{0k}+\theta_{1k}Z_{ik}+e_{Yik},$
@@
where $\theta_{0k}=E\left\{Y_{ik}(0)\mid S_{i}=k\right\};$ $\theta_{1k}=\theta_{Vk}\alpha_{1k}+\gamma_{1k}+\gamma_{2k}\beta_{1k}+\gamma_{3k}\beta_{2k};$ and $e_{Yik}$ has mean zero and is uncorrelated with $Z_{ik}$ within school $k$.
S1.2 School-level population model
Here, we show that $\epsilon_{k}$ defined in model (6) has mean zero and is uncorrelated with the predictors $\beta_{1k},\beta_{2k}$ and $\alpha_{1k}$.